# Proximity and Visuotactile Point Cloud Fusion for Contact Patches in Extreme Deformation

### v1

@@ Abstract
Equipping robots with the sense of touch is critical to emulating the capabilities of humans in real world manipulation tasks. Visuotactile sensors are a popular tactile sensing strategy due to data output compatible with computer vision algorithms and accurate, high resolution estimates of local object geometry. However, these sensors struggle to accommodate high deformations of the sensing surface during object interactions, hindering more informative contact with cm-scale objects frequently encountered in the real world. The soft interfaces of visuotactile sensors are often made of hyperelastic elastomers, which are difficult to simulate quickly and accurately when extremely deformed for tactile information. Additionally, many visuotactile sensors that rely on strict internal light conditions or pattern tracking will fail if the surface is highly deformed. In this work, we propose an algorithm that fuses proximity and visuotactile point clouds for contact patch segmentation that is entirely independent from membrane mechanics. This algorithm exploits the synchronous, high spatial resolution proximity and visuotactile modalities enabled by an extremely deformable, selectively transmissive soft membrane, which uses visible light for visuotactile sensing and infrared light for proximity depth. We present the hardware design, membrane fabrication techniques, and evaluation of our contact patch algorithm in low (10%), medium (60%), and high (100%+) membrane strain states. We compare our algorithm against three baselines: proximity-only, tactile-only, and a first principles membrane mechanics model. Our proposed algorithm outperforms all baselines with an average root-mean-square error (RMSE) under 2.8 mm of the contact patch geometry across all strain ranges. We demonstrate our contact patch algorithm in four applications: varied stiffness membranes, torque and shear-induced wrinkling, closed loop control for whole body manipulation, and pose estimation.
@@ I Introduction
Humans are capable of an exceptionally diverse range of manipulation tasks, effortlessly able to pivot from the precise and delicate motion of threading a needle to the forceful whole-body coordination needed to carry another person.
Figure 1: Illustration of the key idea in fusing proximity and visuotactile point clouds for contact patch segmentation: the set of contact patch points is the intersection of the set of tactile points and the set of proximity points.
The critical role of human skin and sense of touch in the successful execution of these tasks has long been studied with the goal of reproducing human-like manipulation abilities in robots, motivating the rigorous development of electronic skins and tactile sensors [1, 2, 3]. Although there has been much progress, achieving the scalability, coverage, resolution, and compliance of human skin in tactile sensors for robots remains a challenge.
Recently, visuotactile sensors [4, 5, 6, 7, 8, 9] have gained popularity as a tactile sensing strategy due to sensor coverage and resolution that scales with camera field-of-view and pixel count rather than number of sensing elements, as well as compatibility with powerful data processing techniques from computer vision. These sensors use a camera to observe a pattern on a deformable surface, which distorts upon contact. The images of the pattern distortion can inform contact information, such as shear, force, and object geometry.
@@
However, despite leveraging a soft surface as the contact interface, visuotactile sensors rely on sensing mechanisms or data processing techniques that lose functionality beyond small deformations, typically less than 1 mm. For example, dot tracking and optical flow algorithms used to estimate depth or track motion [10, 11, 12, 13] in these sensors are not robust to either occlusions or the large marker displacements that occur during high deformations of soft surfaces. Additionally, photometric stereo and frustrated-internal-reflection methods [14, 15] to estimate depth are heavily dependent on strict lighting conditions that fail when the soft material is highly deformed and too much light is released or the contact area is not evenly illuminated. Because these depth estimation methods limit the depth of object penetration, these sensors use a rigid backplate to prevent further deformation. This creates a critical challenge for visuotactile sensors by fundamentally restricting how much conformal contact can be made with the object.
@@
The lack of conformal contact notably leads to very small contact patches that are not innately useful for pose estimation of cm-scale objects without additional nontrivial implementations of tactile exploration policies [16, 17], extensively training data-driven pose estimation frameworks with experimental or simulated data [18, 19], or integrating external vision sensors in a perception pipeline [20, 21]. While small contact patches may be acceptable for small, mm-scale, distinctly featured objects like screws and nuts, general real-world manipulation demands greater versatility and frequent interactions with larger objects.
During object interactions, limited deformations of the contact interface also increase the difficulty of controlling object pose. With a small range for tolerating shear and torques, objects are much more likely to slip from contact, as interactions with the object closely resemble rigid body interactions [22]. In contrast, a more deformable sensor surface increases the area of stiction with the target object and provides a larger margin of error for robot manipulators through passive compliance, while also allowing a greater range of shear or torque before slip occurs.
@@
Although larger contact patches and deformations are highly desirable, they introduce some new challenges. Because the sensing membrane is compliant, deformation occurs beyond the contact patch with the object. It can be difficult to accurately detect which parts of the membrane are in contact with the object exclusively from sensing the membrane’s deformation. The soft materials used to fabricate these compliant sensing surfaces have highly nonlinear mechanical properties that are difficult to model accurately and compute in real time for larger deformations.
@@
In this work, we propose a contact patch segmentation algorithm that fuses proximity point clouds and tactile point clouds to detect contact patches of objects in extreme deformations of the sensing membrane. This algorithm leverages the multimodality of a large-scale proximity and visuotactile sensor enabled by a novel selectively transmissive soft membrane. The proximity depth modality is used to enhance tactile sensing functionality by providing essential object localization and geometry data during contact that is independent from the mechanical behavior of the soft membrane. With this key information and collocated proximity and tactile depth sensors, the intersection of the proximity and tactile point clouds provides the contact patch (Figure 1). We make three key contributions:
@@
    We present and evaluate our contact patch algorithm across low, medium, and high strain membrane states and against three baselines: proximity-only, tactile-only, and a membrane mechanics model. To the best of our knowledge, this is the first contact patch segmentation algorithm that fuses a point cloud proximity modality with tactile sensing.
@@
    We introduce design and fabrication techniques for a multimodal proximity and visuotactile sensor enabled by a selectively transmissive soft membrane.
  3. 3.
@@ II Related Work
II-A Multimodal Proximity and Tactile Sensors
Multimodal proximity and tactile sensors have been primarily explored in the form of sensing skins or sensing modules covering a robot end effector [23]. These sensing skins and modules commonly use a MEMs-based approach to measure proximity, integrating time-of-flight distance sensors to provide single point measurements for pre-grasp object detection [24, 25, 26] or re-grasping reflexes [27, 28]. However, the form factor of these individual MEMs proximity sensors are bulky and have wiring requirements that do not scale well for higher spatial resolution sensing. The lack of higher spatial resolution tactile and proximity sensing from the MEMs chips approach can severely hinder higher-level perception tasks such as object recognition, object tracking, and contact patch estimation.
FingerVision [29] is one of the few proximity and visuotactile sensors that provide a proximity modality, by using stereo vision to observe a pattern on a clear elastomer. However, this design requires a fundamental compromise in the spatial resolutions of proximity and tactile modalities: adding denser features to the pattern for tactile sensing occludes the proximity vision, while a sparse pattern limits the tactile resolution.
See-Through-Your-Skin [30, 31] is another proximity and visuotactile sensor that provides a proximity modality by modulating the membrane transparency via the two-way mirror effect with a semi-reflective surface. However, only one modality can be active at a time due to the transparency modulation: either proximity or tactile, but not both. Additionally, the two-way mirror sensing mechanism is not functional in high deformations, as the semi-reflective surface lets in too much external light and the sensor would be limited to the proximity modality.
Protac [32] is a proximity and visuotactile sensor that uses an electrically-controlled transparency-modulated film to switch between the tactile and proximity modalities. This approach, however, faces limitations in time resolution: the modalities are not synchronous and it takes 0.3s to switch between the modalities, which is a significant delay for real-time object interactions.
In this work, we overcome these limitations by using a selectively transmissive soft membrane [33], which blocks visible light and allows infrared light to pass through. We use an infrared time-of-flight depth camera to see through the membrane for the proximity modality, and a stereo RGB camera to sense the membrane for the tactile modality. Because the modalities do not interfere with each other, we achieve time-synchronous and simultaneous proximity and tactile point clouds.
II-B Contact Patch Estimation
@@
Contact patch estimation is an important functionality for tactile sensors, often the first step in processing tactile data for downstream tasks like pose estimation, estimating contact force distributions and direction, and object classification. The three main approaches to contact patch estimation for soft tactile sensors with hyperelastic materials are to apply a threshold, use data-driven techniques to predict the contact patch, or solve a finite element method (FEM) model online.
@@
Applying a threshold to displacement values to estimate the contact patch is the simplest and fastest method [22]. Typically a hand-selected number is used as the threshold and membrane displacement beyond the threshold is considered as part of the contact patch estimate. However, thresholds require manual tuning and do not generalize well across a large range of deformations.
@@
Machine learning methods have been proposed to predict the contact patch but this approach brings another set of significant challenges: (1) the ability to generalize to a diverse set of object geometries, and (2) the laborious process of generating experimental datasets and an associated ground truth. Because the displacement to force mapping is dependent on object geometry, the datasets would have to be overwhelmingly extensive to capture a fraction of possible objects that could be encountered by the sensor. Additionally, the ground truth for each image would have to be generated by a corresponding FEM simulation [34, 35], for which accuracy scales with computation cost.
@@
With the FEM approach, each element of the elastomer is typically modelled as a hyperelastic material, for which the stress-strain relationship is derived from a strain energy density function [36]. Commonly used hyperelastic models for elastomers include the Neo-Hookean, Ogden, and Mooney-Rivlin formulations. However, to determine the stress-stretch curves necessary to use these models, samples of the material must be tested in uniaxial tension, pure shear, and equibiaxial tension [37]. Collecting this type of data requires specialized equipment and does not generalize well across manufacturing variabilities in different samples. Most significantly, FEM approaches are very computationally expensive and do not run in real-time. Linear approximations or assumptions of a constant stiffness matrix can be made to speed up runtime [38, 39], but these assumptions do not hold in cases of extreme deformations.
III Sensor System Overview
The design and fabrication of the sensor builds upon [33]. The sensor provides synchronized 640x480 px depth maps in both the proximity and tactile modalities at 30 Hz. The soft sensing surface is 355 mm x 205 mm and allows up to 85 mm of vertical displacement from a deflated membrane, before reaching the minimum sensing distances of the internal cameras. The sensor can be pressurized to inflate the membrane; the inflated height directly corresponds to how much the membrane can be vertically displaced. The proximity depth range extends approximately 100 mm beyond the membrane.
Figure 2: Photos of the selectively transmissive soft membrane (top) and integration with the sensor system (middle and bottom).
III-A Selectively Transmissive Soft Membrane
We designed the new membrane presented in this work to achieve the following:
  1. 1.
    Provide a rich set of dense and random features in visible wavelengths (400 nm - 700 nm) to the internal stereo RGB camera for tactile point clouds.
  2. 2.
    Allow the infrared light (860 nm) emitted by the internal time-of-flight camera to pass through for the proximity point cloud.
  3. 3.
    Be produced from a simple, nontoxic fabrication process that facilitates easy repairs.
Figure 3: The cup with a handle is pressed into the sensing surface of the multimodal visuotactile and proximity sensor by a robot arm. The corresponding sensing data is shown in the middle, where the contact patch estimation is computed by fusing the proximity and tactile point clouds. The ground truth contact patch is shown on the right.
The soft membrane consists of two layers of silicone elastomer. The first layer is fully cured prior to pouring the next layer directly on top. The base elastomer is Ecoflex 00-35 Fast (Smooth-On), which we observe to have similar optical properties to Ecoflex 00-30 (Smooth-On) that has an attenuation of approximately 10db/cm at 860nm [40]. With the final thickness of the membrane at 1 mm, the membrane transmits approximately 90% of the emitted infrared (IR) light from the depth camera. In case of any small punctures or rips, the membrane can be quickly repaired by pouring another thin layer of the same silicone directly at the location of damage with no impact to sensing functionality.
The first layer consists of an IR-transparent, visibly-colored aqueous dye solution (VIS691A; QCR Solutions Corp, 0.1g/100mL concentration) hand-stirred with Ecoflex 00-35 Fast at a ratio of 1:10 (mL) dye to elastomer ratio. For the membrane in this work, we use 100 g of elastomer and 10 mL of dye solution. The dye solution is insoluble in silicone, which enables the formation of randomly encapsulated dye droplets to create the necessary visual pattern for depth estimation by the stereo RGB camera. The size of the dye droplets can be partially controlled by the amount of time spent mixing the dye and elastomer together; more time spent mixing is correlated with smaller dye droplets. The desired size of the dye droplets is around 1-3 mm; this is large enough for the RGB stereo camera to detect, small enough for inconsequential stress concentrations introduced into the silicone elastomer, and creates a sufficient density of droplets for a rich visual pattern. Then, we mix 35 g of NOVOCS Matte (Smooth-On) with the dye and elastomer mixture. NOVOCS Matte imparts a matte finish on the membrane to reduce internal specular reflections and thins the elastomer mixture for even pouring into the mold. We pour the dyed elastomer into a laser-cut acrylic mold and let it fully cure at room temperature, which takes 5 minutes.
The final layer makes the membrane airtight by sealing any small pores in the membrane created by the encapsulated dye droplets, since the membrane will be inflated for use. We mix 30 g of Ecoflex 00-35 Fast with 20 g of NOVOCS Matte and pour it onto the first layer. The layer is then fully cured at room temperature in 5 minutes. For increased physical robustness, this final sealing layer can be increased in thickness at the cost of a decreased range of proximity sensing. For an increased range in proximity sensing, the silicone can be poured selectively (rather than an entire layer) on potential air leaks, which can be inferred by the locations of larger dye droplets.
III-B Internal Electronics
@@
The Intel Realsense L515, which has integrated RGB and time-of-flight (ToF) cameras, is used for proximity sensing. The IR light emitted by the L515 ToF camera travels through the membrane and does not sense the IR-transparent membrane itself. The Intel Realsense D405, a stereo RGB depth camera, is used for tactile sensing. Although the membrane is somewhat translucent, the visual texture of the membrane causes the D405 to produce a point cloud of the membrane only. The two cameras are placed immediately next to each other at the bottom center of the sensor. We design the active sensing region to be located where the FOVs of the D405’s stereo depth and L515’s ToF camera overlap; the entire exposed membrane is in view of both Realsense cameras. The cameras output data through USB-C ports, which are connected to an airtight USB-A 3.0 port that goes through the sensor housing. Both Realsense cameras are sampled synchronously at 30 FPS, which is the fastest frame rate from the Realsense L515 ToF camera.
To provide internal lighting for the Realsense D405, we mount two LED ring lights to the sides of the sensor housing. Additionally, we include an internal air pressure sensor and a microcontroller that are also connected to the airtight USB-A port for data transmission. The internal air pressure sensor is required for the membrane mechanics model baseline comparison discussed in Section VC. An overview of the sensor system is shown in Figure 2.
III-C Sensor Packaging
The sensor in this work is designed for integration into a whole-body robot manipulation system, so the sensing surface is scaled to cover the torso of the robot. The sensor can also be scaled down to be used as an end effector of a robot arm [33], though the size is limited by the minimum sensing distance of the L515’s depth camera. The sensor housing is 3D printed (VeroBlack, J750, Stratasys) and consists of the main sensor chamber (403 mm x 280 mm x 284 mm) and lid (425 mm x 276 mm x 44 mm). The sensor chamber has rails compatible with 80/20 components for mounting to the robot body, as well as a push-to-connect tube fitting and two USB-A 3.0 ports. The lid features slots for M3 nuts and the chamber is attached with M3 screws. The lid clamps down on the membrane, sandwiched by two silicone gaskets, to form an airtight seal.
@@ IV Proximity and Visuotactile Fusion for Contact Patch Estimation
The fusion algorithm outputs a point cloud of the estimated contact patch, which is constructed from the intersection of the set of points from the tactile and proximity point clouds shown in Figure 1 and Figure 3.
The data inputs to the algorithm are the proximity camera’s RGB image, the proximity depth image, and the tactile depth image. The proximity camera’s RGB image, which sees the membrane, is only used to remove noise from the proximity depth data. The proximity depth and tactile depth maps are used to compute the contact patch.
@@ IV-2 Proximity Pre-Processing
Depending on the curvature of the inflated membrane, some areas of the membrane can be infrared reflective and produce noise in the proximity depth data. We convert the proximity RGB image to HSV and use manually tuned thresholds to create a mask that removes any pixels that match the color of the membrane when no objects are in contact or near the membrane surface. The mask is then applied to the proximity depth image, which filters out the internal reflections of the emitted infrared light.
@@ IV-3 Contact Patch Estimation
where $d_{p}$ is the proximity distance, $d_{t}$ is the tactile distance, and $t$ is the tolerance. If the difference in tactile and proximity distance values at a pixel satisfy the inequality, the pixel is considered a match and part of the contact patch. We manually tune the value of $t$ to produce the contact patch that best matches ground truth, and we found that $t=0.03$ works well; in our experiments, the tolerance distance values range from 6mm-8mm. The intersection mask is applied to the proximity depth image to produce the contact patch estimation. We use the proximity depth image instead of the tactile depth image because the proximity depth data is significantly more robust to inconsistent room lighting conditions. The contact patch depth image is then projected to a point cloud and we apply a statistical outlier rejection filter to remove any extraneous points before outputting the final contact patch point cloud estimation.

### v2

@@ Abstract
Visuotactile sensors are a popular tactile sensing strategy due to high-fidelity estimates of local object geometry. However, existing algorithms for processing raw sensor inputs to useful intermediate signals such as contact patches struggle in high-deformation regimes. This is due to physical constraints imposed by sensor hardware and small-deformation assumptions used by mechanics-based models. In this work, we propose a fusion algorithm for proximity and visuotactile point clouds for contact patch segmentation, entirely independent from membrane mechanics. This algorithm exploits the synchronous, high spatial resolution proximity and visuotactile modalities enabled by an extremely deformable, selectively transmissive soft membrane, which uses visible light for visuotactile sensing and infrared light for proximity depth. We evaluate our contact patch algorithm in low (10%), medium (60%), and high (100%+) strain states. We compare our method against three baselines: proximity-only, tactile-only, and a first principles mechanics model. Our approach outperforms all baselines with an average RMSE under 2.8 mm of the contact patch geometry across all strain ranges. We demonstrate our contact patch algorithm in four applications: varied stiffness membranes, torque and shear-induced wrinkling, closed loop control, and pose estimation.
@@ I Introduction
Humans are capable of an exceptionally diverse range of manipulation tasks, from precisely threading a needle to carrying another person. Tactile sensing has been studied with the goal of reproducing human-like manipulation abilities in robots. A popular tactile sensor paradigm are visuotactile sensors, which utilize high-resolution cameras to leverage advancements in computer vision techniques [1, 2, 3, 4, 5, 6].
Figure 1: We propose a mechanics-independent algorithm for estimating contact patches in challenging deformation regimes. We leverage a selectively transmissive soft membrane to provide simultaneous tactile and proximity point clouds. By fusing these two modalities through simply computing the intersection between the two point clouds, we obtain a high quality estimate of the contact patch. Because of our mechanics-independent approach, we avoid the non-linearities that many existing contact patches algorithm struggle with.
@@
However, despite leveraging a soft surface as the contact interface, most visuotactile sensors rely on data processing techniques that lose functionality beyond small deformations, typically less than 1 mm. For example, dot tracking and optical flow algorithms used to estimate depth or track motion [7, 8, 9, 10] in these sensors are not robust to either occlusions or the large marker displacements that occur during high deformation. This fundamentally restricts how much conformal contact can be made with the object.
@@
The lack of conformal contact notably leads to small contact patches that are not innately informative for cm-scale objects without additional nontrivial implementations of tactile exploration policies [11, 12], data-driven pose estimation frameworks [13, 14], or integrating external vision sensors [15, 16]. While small contact patches may be acceptable for small, mm-scale, distinctly featured objects like screws and nuts, general real-world manipulation demands greater versatility and frequent interactions with larger objects. Limited conformal contact can also increase the difficulty of controlling the object. More compliance can be beneficial as it increases stiction and provides a larger margin of error for robots [17]. However, existing algorithms are poorly equipped to operate in the these regimes of large compliance.
@@
Because deformation occurs beyond the contact patch, it is difficult to detect which parts of the membrane are in contact with the object exclusively from sensing the membrane’s deformation. The soft materials used to fabricate these compliant sensing surfaces have highly nonlinear mechanical properties that are difficult to model accurately and compute in real time for larger deformations.
@@
In this work, we propose a contact patch segmentation algorithm that fuses proximity point clouds and tactile point clouds to detect contact patches of objects in extreme deformations of the sensing membrane. This algorithm leverages the multimodality of a large-scale proximity and visuotactile sensor enabled by a selectively transmissive soft membrane. The proximity depth modality is used to enhance tactile sensing functionality by providing essential object localization and geometry data during contact, independent from the mechanical behavior of the soft membrane. With this key information and collocated proximity and tactile depth sensors, the intersection of the proximity and tactile point clouds provides the contact patch (Fig. 1). We make two key contributions:
@@
    We propose a novel contact patch segmentation algorithm that is independent from mechanics of the deformable membrane. This simple algorithm enables robustness to challenging deformation states, while still being fast enough for real-time control feedback. To the best of our knowledge, this is the first contact patch segmentation algorithm that fuses a point cloud proximity modality with tactile sensing.
@@ II Related Work
Contact patch estimation is often the first step in processing tactile data for downstream tasks such as estimating object pose, contact force distributions and direction, and object classification. The three main approaches to contact patch estimation for soft tactile sensors with hyperelastic materials are to apply a threshold based on deformation, use data-driven techniques to classify contact patches, or solve a finite element method (FEM) model online.
@@
Applying a threshold to membrane displacement is the simplest and fastest method to estimate contact patches [17]. A membrane displacement threshold number is manually chosen, and points over the threshold are considered as the contact patch. However, thresholds require manual tuning specific to the expected deformation and contact interface material, and do not generalize well across a large range of deformations.
@@
Machine learning methods have been used to predict the contact patch, but they still have significant challenges: (1) the ability to generalize to a diverse set of object geometries, and (2) the laborious process of collecting real-world labelled datasets. Because the displacement to force mapping is dependent on object geometry and membrane mechanics, the datasets have to be very extensive to generalize to all possible objects that could be encountered by the sensor. Additionally, the ground truth for each image would have to be generated by a corresponding FEM simulation [18, 19], for which accuracy scales with computation cost.
@@
With the FEM approach, each element of the elastomer is modelled as a hyperelastic material, for which the stress-strain relationship is derived from a strain energy density function [20]. Commonly used hyperelastic models for elastomers include the Neo-Hookean, Ogden, and Mooney-Rivlin formulations. However, to determine the stress-stretch curves necessary to use these models, samples of the material must be tested in uniaxial tension, pure shear, and equibiaxial tension [21]. Collecting this data requires specialized equipment and does not generalize well across manufacturing variabilities in different samples. Most significantly, FEM approaches are computationally expensive and do not run in real-time. Linear approximations or assumptions of a constant stiffness matrix can be made to speed up runtime [22, 23], but these assumptions do not hold in cases of extreme deformations.
III Sensor System
The design and fabrication of the sensor builds upon [24]. The sensor provides synchronized 640x480 px depth maps in both the proximity and tactile modalities at 30 Hz.
@@ III Sensor System
The two internal cameras are the Intel Realsense L515 Time-of-Flight camera (proximity modality) and the Intel Realsense D405 (tactile modality). The IR light emitted by the L515 ToF camera travels through the membrane and does not sense the IR-transparent membrane itself. Although the membrane is somewhat translucent, the visual texture embedded in the membrane with random dye droplets enables the D405 to produce a point cloud of the membrane. The two cameras are placed immediately next to each other at the bottom center of the sensor. We design the active sensing region to be located where the FOVs of the D405’s stereo depth and L515’s ToF camera overlap; the entire exposed membrane is in view of both cameras.
The soft sensing surface is 355 mm x 205 mm and allows up to 85 mm of vertical displacement from a deflated membrane, before reaching the minimum sensing distances of the internal cameras. The sensor can be pressurized to inflate the membrane; the inflated height directly corresponds to how much the membrane can be vertically displaced. The proximity depth range extends approximately 100 mm beyond the membrane.
@@ IV Proximity and Visuotactile Fusion for Contact Patch Estimation
The fusion algorithm outputs a point cloud of the estimated contact patch, which is constructed from the intersection of the tactile and proximity point clouds (Fig. 1).
The data inputs to the algorithm are the proximity camera’s RGB image, the proximity depth image, and the tactile depth image. The proximity camera’s RGB modality, which observes the membrane, is only used to remove noise from the proximity depth data. The proximity depth and tactile depth maps are used to compute the contact patch.
@@ IV-2 Proximity Pre-Processing
Depending on the curvature of the inflated membrane, some areas of the membrane can be infrared reflective and produce noise in the proximity depth data. We convert the proximity RGB image to HSV and create a mask that removes any pixels that match the color of the membrane when no objects are in contact or near the membrane surface. This mask is tuned to environmental lighting conditions, but applies to all objects. The mask is then applied to the proximity depth image, which filters out the internal reflections of the emitted infrared light.
@@ IV-3 Contact Patch Estimation
where $d_{p}$ is the proximity distance, $d_{t}$ is the tactile distance, and $t$ is the tolerance. If the difference in tactile and proximity distance values at a pixel satisfy the inequality, the pixel is considered a match and part of the contact patch. We manually tune the value of $t$ to produce the contact patch that best matches ground truth, and we found that $t=0.03$ works well; the tolerance distance values range from 6mm-8mm. The intersection mask is applied to the proximity depth image to produce the contact patch estimation. We use the proximity depth image instead of the tactile depth image because it is significantly more robust to inconsistent room lighting conditions. The contact patch depth image is then projected to a point cloud and we apply a statistical outlier rejection filter to remove any extraneous points before outputting the final contact patch point cloud estimation. Although this step does require some manual tuning of $t$, it generalizes across all strain ranges and objects.